OpenStudy (anonymous):

18x - 7 = 10x + 33 *

OpenStudy (anonymous):

Please help me im confused...

OpenStudy (welshfella):

i guess you want to solve this for x

OpenStudy (anonymous):

yeah

OpenStudy (welshfella):

we aim to get all the terms in x on left side and numbers on the right 18x - 7 = 10x + 33 first subtract 10x from both sides of the '=' 18x - 10x - 7 = 10x - 10x + 33 simplifying; 8x - 7 = 33 Now add 7 to both sides:- 8x - 7 + 7 = 33 + 7 8x = 40 finally to get x , divide both sides by 8 8x 40 -- = -- 8 8 x = 5

OpenStudy (welshfella):

so x = 5 is the solution

OpenStudy (anonymous):

ok thank you

OpenStudy (welshfella):

yw